OpenStudy (anonymous):

For two events M and N, P(M)=0.6, P(N|M)=0.3, P(N|M')=0.2. Find P(M|N')

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

i said hi

OpenStudy (anonymous):

hi! @Gabbar

OpenStudy (anonymous):

the second one should be P(N'|M) =0.3, right?

OpenStudy (anonymous):

If it is so, using Baye's theorem to find P(M|N') If it is not, I don't know how to solve. :)

OpenStudy (anonymous):

no thats what's confusing me is why it's set up that way

OpenStudy (anonymous):

unless its a typo

OpenStudy (anonymous):

hihihi, I am useless then. I am sorry. @ganeshie8

OpenStudy (xapproachesinfinity):

is the N' the complement of N?

OpenStudy (xapproachesinfinity):

Another prob problem hsha. you sure like this aren't you? hehe but seriously probs are pretty amazing stuff

OpenStudy (anonymous):

haha im not having too much fun with them

OpenStudy (xapproachesinfinity):

Okay! you should be having fun hehe

OpenStudy (xapproachesinfinity):

okay let's give a shot! P(M/N')=P(N'/M)P(M)/P(N'/M)P(M)+P(N'/M')P(M') is this how we wrote that thing before?

OpenStudy (xapproachesinfinity):

well this should the thm that oops talked about!

OpenStudy (anonymous):

i think it may have been a typo but ill work that out!

OpenStudy (xapproachesinfinity):

So far i got \(\large\frac{P(N' and M)}{P(N' and M)+0.4\times0.7}\)

OpenStudy (xapproachesinfinity):

we need to find that P(N' and M) to get this hehe

OpenStudy (xapproachesinfinity):

I guess there is not typo you just have to find a way through this!

OpenStudy (xapproachesinfinity):

Any Ideas! more minds are always better hehe. i might be missing something that you guys paid attention to

OpenStudy (xapproachesinfinity):

i based this if N' is the complement of N

OpenStudy (anonymous):

so N'=0.7 right?

OpenStudy (xapproachesinfinity):

where did you got that? 0.7 is complement of P(N' given M') not P(N') apparently we don't have P(N) so no P(N') either hehe

OpenStudy (anonymous):

ahh nevermind i was getting it from P(N|M)

OpenStudy (xapproachesinfinity):

well that's the complement of P(N' given M') that's why it 0.7

OpenStudy (xapproachesinfinity):

"0.7 is complement of P(N' given M') not P(N')" Actually what i said here is not true it is the opposite i meant the complement for P(N/M)

OpenStudy (xapproachesinfinity):

Anyways we need something to get that P(N and M')

OpenStudy (anonymous):

we need to find P(N) right?

OpenStudy (xapproachesinfinity):

I'm looking for P(n) to get P(n') we want to get P(N' and M) correction for the line i made earlier

OpenStudy (xapproachesinfinity):

Apparently we have P(N/M')=0.2

OpenStudy (xapproachesinfinity):

so P(N/M')=P(N)P(M')/P(M')=0.2 hence P(N)=0.2 So P(N')=1-0.2=0.8

OpenStudy (xapproachesinfinity):

does that sound fine?

OpenStudy (xapproachesinfinity):

we already have P(M) so P(N' and M)=P(N')P(M)=0.8\(\times\)0.6=0.48

OpenStudy (xapproachesinfinity):

\(\huge \frac{0.48}{0.48+0.4\times0.7}\)

OpenStudy (xapproachesinfinity):

I'm not sure im doing this good hehe. i might be used some wrong steps xD

OpenStudy (xapproachesinfinity):

ganesh is this P(N'/M') complement of P(N/M) not sure that's true!

OpenStudy (xapproachesinfinity):

Actually we still have to ask are we dealing with independent events or not?

ganeshie8 (ganeshie8):

yes bayee's theorem works :) we may use contingency table : 

OpenStudy (xapproachesinfinity):

Oh that pretty good, never used such table ^_^

ganeshie8 (ganeshie8):

filling in the given info : ` P(M)=0.6, P(N|M)=0.3, P(N|M')=0.2.`